What is the pH of a buffer made from 0.580 mol of CH₃NH₂ (Kb = 4.4 × 10⁻⁴) and 0.220 mol of CH₃NH₃I?
Answer #1

EMA] рои - речь ) уоu - pk t log [ си, чи, ) ги, и, 2] pkb = -log (kb) Кь = -0, 14-Чx IsЧ) Рkb = 3- 35 рои = 3-35 лоу | | ро